1. why is 9 a perfect square?
2. represent 125 in exponental from?

Answer by mananth(16949)

the square of a number is one number mutiplied by itself. This number has to be rational.
3*3=9 so 9 is a perfect square
...
the factors of 125 are 5*5*5
=5^3
5^3=125
...
m.ananth@hotmail.ca